OpenStudy (anonymous):

What transformations change the graph of f(x) to the graph of g(x).... F(x) =2x. G(x)=6x-3

OpenStudy (chrisplusian):

Is that how the problem is worded?

OpenStudy (chrisplusian):

Is vertical stretching and a vertical shift.

OpenStudy (anonymous):

Apparently there is a problem w the equation. It gives these answers.... G(x) =6x-3 =3(2x)-3 =3f(x)-3 I don't know where the error is

OpenStudy (chrisplusian):

If I had to guess at one I would say it is =3f(x)-3 because they have taken the graph of "f(x)" and vertically stretched it by a factor of three and transposed it three places down so I would guess at that one.

OpenStudy (anonymous):

That is what i have been saying this whole time too, but what is the problem with it i cant locate it ?

OpenStudy (chrisplusian):

Is that not an option? If not f(x)=2x so technically you could say that it is also 3(2x)-3

OpenStudy (chrisplusian):

did that work?

OpenStudy (anonymous):

i believe that the last portion should'nt even be in there i very much appreciate all the help!

OpenStudy (chrisplusian):

No problem :)